A certain upgrade causes a high-tech space probe to travel at 5 light years per century instead of 3.75 light years per century, so that it can reach the Horsehead Nebula 10 000 years faster. How far away is the Horsehead Nebula, in light years?
Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

A certain upgrade causes a high-tech space probe to travel at 5 light years per century instead of 3.75 light years per century, so that it can reach the Horsehead Nebula 10 000 years faster. How far away is the Horsehead Nebula, in light years?
-----
Original DATA::
dist to HN = x light years ; rate = 3.75 ly/c ; time = x/3.75 c
----
Changed DATA:
dist to HN = x ly ; rate = 5 ly/c ; time = x/5 c
-------------
Equation:: x/3.75 c - x/5 c = 100 c
-------------------------------------
5x - 3.75x = 3.75*5*100
-----
1.25x = 5*375
Answer:: x = 5*3 = 15 light years
-----------
Cheers,
Stan H.
-----------

Answer by ikleyn(52776)

.
The correct answer is 1500 light years.

Compare it with this Wikipedia article

    https://en.wikipedia.org/wiki/Horsehead_Nebula